use the five steps for problem solving to answer the following question. please show all of your work.
the average of two test scores is 79 if one test score is eleven less than the other, what are the two test scores?

Answer by heilok(4)

Hi.

Let x be the score of the test with better performance.
Then the score of another test is (x-11).

Average score of the 2 tests=79
[x+(x-11)]/2=79
2x-11=158
x=84.5

The score of the test with better result is 84.5
The score of another test=84.5-11=73.5.

Hope it helps :)